OpenStudy (rainbow_rocks03):

PLZ HELP ME I will medal you for the best response.Solve for g. g – (–43) = 19 A. –62 B. –24 C. 24 D. 62

OpenStudy (idealist10):

g-(-43)=19 g+43=19 g=19-43 g=-24 Answer: -24, B.

OpenStudy (anonymous):

g-(-43)=19 => g+43=19

OpenStudy (anonymous):

Now just subtract 43 on both sides

OpenStudy (rainbow_rocks03):

Ok I thought it was B.

OpenStudy (rainbow_rocks03):

so -24-43 @iambatman ?

OpenStudy (anonymous):

Where did you get the -24

OpenStudy (anonymous):

its 19-43

OpenStudy (rainbow_rocks03):

oh so 19- -43?

OpenStudy (anonymous):

I don't know why you're doing what you're doing.

OpenStudy (rainbow_rocks03):

So the answer is D @iambatman ?

OpenStudy (idealist10):

When you subtract a negative number, you just add that number.

OpenStudy (anonymous):

When you have a problem as yours, the point is to get the variable (in your case g) on its own

OpenStudy (anonymous):

To do this, you basically do the opposite operation of what already exists to move a number to the other side.

OpenStudy (anonymous):

g - (-45) = 19 negative times a negative = positive g+45=19 -45 -45 you see 45's get cancelled on the left and you must do what you do on one side to the other. so g = 19-45

OpenStudy (rainbow_rocks03):

So it is B?

OpenStudy (rainbow_rocks03):

I got it right it was B.

OpenStudy (idealist10):

Told you it's B. :)